OpenStudy (anonymous):

HELP PLEASE The volume of a cube is 54 cubic feet. What is the volume of a pyramid that fits exactly inside the cube?

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

@GretaKnows

OpenStudy (anonymous):

@Godlovesme

OpenStudy (godlovesme):

no idea sowwy :/

OpenStudy (anonymous):

same idk :(

TheSmartOne (thesmartone):

Volume of a cube = \(\sf side^3\) Volume of a pyramid: \(\sf\Large \frac{length \times width \times height}{3}\)

TheSmartOne (thesmartone):

They gave us the volume of the cube is 54. Plug that into the first formula and solve for side. And then since the pyramid can not be bigger than the length of the side of a cube because it has to fit inside that means that the length, width and height are equal to the side. So the formula of the volume will basically be: \(\sf\Large\frac{side^3}{3}\)

TheSmartOne (thesmartone):

But since we already know what \(\sf side^3\) equals to, the answer becomes simpler. So your answer is \(\sf\Large\frac{side^3}{3}\) So plug in \(\sf side^3 = 54\) and simplify it :)